OpenStudy (anonymous):

what is X when 2sin^2(X)+9cos(X)+3=0

OpenStudy (anonymous):

have you subbed \(\sin^2 x x = 1-\cos^2 x\) into the eq to simplify it into a quadratic eq?

OpenStudy (anonymous):

no i haven't but i'll try it out right now thanks :)

OpenStudy (anonymous):

u r welcome. The solution should be trivial after the substitution. :)

OpenStudy (anonymous):

I tried it but i still don't get it..

OpenStudy (anonymous):

subbing, you should get, \(2-2\cos^2 x + 9 \cos x +3=0\) which equals, \(-2\cos^2 x + 9 \cos x +5=0\) let u = cos x the equation becomes \(-2u^2 + 9 u +5=0\) which is a quadratic equation of u. You can use the general formula to solve for u.

OpenStudy (anonymous):

thanks o much.. i really appreciate it

OpenStudy (anonymous):

you're welcome :)